[litho] Morning Litho Status 4/24
Gary J Sosa
gsosa at stanford.edu
Wed Apr 24 07:49:28 PDT 2013
Good morning All..
General Information/Lab Activities
Wednesday, April 24, 2013
Intermittent BSA focus issue
Making arrangements for annual PM . PO Approved. Scheduling of PM’s for both aligners in progress
Does not work as standard aligner. Suspect software or configuration
Further investigation shows that system does not work in imprint mode either. Network and internet functional. Attempting an online session with EVG this morning to diagnose system
Bond-Align not working
PO for service approved on 4/10. Working with EVG on comprehensive plan to check out/repair his tool on the imprint tool.
Cannot change temperature
Arranging with supplier for new temperature controller
Continuous testing and investigating
1.Chillplate temp control
2. Wafer Handling
1. Replaced water pump. Testing in progress.
2. Testing various paths and wafer sizes.
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the litho